OpenStudy (anonymous):

Which statement describes the graph of the equation x^2+y^2+4x-6y-3=0? A) A hyperbola with center (-2,3) and vertices (4,-3) and (-4,3) B) a hyperbola with center (-2, 3) and vertices (2,-3) and (3,-2) C) a circle with center (-2,3) and radius 8 D) a circle with center (-2,3) and radius 4

OpenStudy (anonymous):

@sparky16 try to write the equation as (x^2+4x) + (y^2-6y) -3 = 0 now can u make the perfect squares?

OpenStudy (anonymous):

OR can u write x^2 + 4x in the form of (x + some constant value)^2 ???????

OpenStudy (anonymous):

so I got down to this (x+2)^2+(y-3)^2=11 where do I go from here?

OpenStudy (anonymous):

u r going to get (x+2)^2 + (y-3)^2 = 16

OpenStudy (anonymous):

oh yes your right it equals 16

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

the result we get is equation of circle equation of circle is (x-a)^2 + (y-b)^2 = r^2 here 'r' is the radius of the circle and 'a' and 'b' are coordinates of center.

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

yup